OpenStudy (anonymous):

Express the complex number in trigonometric form. -3i

hartnn (hartnn):

The formula to convert x+iy into r cos theta + i r sin theta is : r = \(\sqrt{x^2+y^2}\) \(\theta = \tan^{-1}(\dfrac{y}{x})\)

hartnn (hartnn):

so, in your case, -3i = \(\large 0-3i\) x= 0, y = -3 plug these in the formula.

OpenStudy (anonymous):

so r=3 i tried the 2nd formula but i got an error on my calculator

hartnn (hartnn):

right, thats because tan 90 is infinity which the calculator does not support displaying. use the fact that tan inverse (negative number/ 0) = - 90 so theta = -90

OpenStudy (anonymous):

ok so now what

hartnn (hartnn):

r and theta define your trigonometric form, r angle theta -3i = 3 angle -90 thats it!

OpenStudy (anonymous):

but my choices are different compared to your answer. A) 3(cos 180° + i sin 180°) B) 3(cos 270° + i sin 270°) C) 3(cos 90° + i sin 90°) D) 3(cos 0° + i sin 0°)

hartnn (hartnn):

ok, so you're clear with r being 3, right ?

OpenStudy (anonymous):

yup

hartnn (hartnn):

now we need -i part, which means cos value should be 0 cos is 0 when angle is 90 or 270

hartnn (hartnn):

and since we need sin part as -1 we know sin 270 = -1

hartnn (hartnn):

so, finally, 3(cos 270° + i sin 270°)

OpenStudy (anonymous):

oh i get, thank you so much!

hartnn (hartnn):

welcome ^_^